### v2

# Beyond Multi-Token Prediction: Pretraining LLMs with Future Summaries

Abstract
Next-token prediction (NTP) has driven the success of large language models (LLMs), but it struggles with long-horizon reasoning, planning, and creative writing, with these limitations largely attributed to teacher-forced training. Multi-token prediction (MTP) partially mitigates these issues by predicting several future tokens at once, but it mostly captures short-range dependencies and offers limited improvement. We propose future summary prediction (FSP), which trains an auxiliary head to predict a compact representation of the long-term future, preserving information relevant for long-form generations. We explore two variants of FSP: handcrafted summaries, for example, a bag of words summary of the future sequence, and learned summaries, which use embeddings produced by a reverse language model trained from right-to-left order. Large-scale pretraining experiments (3B and 8B-parameter models) demonstrate that FSP provides improvements over both NTP and MTP across math, reasoning, and coding benchmarks.
[…]
1 Introduction
Early progress in large language models was primarily driven by massive scaling of both data and compute (Brown et al., 2020; Kaplan et al., 2020). However, the returns from this vanilla scaling approach are beginning to diminish as we encounter the “data wall” (Villalobos et al., 2024; Sutskever, 2024). This has renewed efforts toward algorithmic advances, including new architectures and pretraining objectives, that can extract more predictive signal from a fixed amount of training data (Muennighoff et al., 2023).
Next-token prediction (NTP) with teacher forcing—training models by conditioning on ground-truth history when predicting the next token—is foundational to current pretraining methods. However, this approach introduces a train-inference mismatch known as exposure bias: during inference, the model must rely on its own outputs rather than the ground truth, leading to compounding errors and degraded long-range generation quality (Bengio et al., 2015; Ranzato et al., 2015). Moreover, teacher forcing can induce training-time shortcut learning as well, where the model exploits local cues from the ground-truth prefix instead of capturing true long-range dependencies (Bachmann and Nagarajan, 2024). These issues manifest most clearly in tasks demanding long-horizon planning, narrative coherence, and open-ended creativity (Papalampidi et al., 2022; Nagarajan et al., 2025).
An appealing alternative to teacher forcing is teacherless training, where the model learns from its own generations rather than relying on ground-truth histories. However, this approach is computationally intensive and challenging to parallelize. As a practical compromise, recent research has explored multi-token prediction (MTP) methods (Gloeckle et al., 2024), which train auxiliary heads to predict several future tokens simultaneously. MTP has demonstrated success in large-scale systems such as DeepSeek-V3 (Liu et al., 2024), Qwen-3 (Yang et al., 2025), MiMo (Xiaomi et al., 2025), and Nemotron-3 (Blakeman et al., 2025). In MTP, each time step augments the standard next-token prediction ($x_{t+1}$) with auxiliary heads that predict additional future tokens (e.g., $x_{t+2}$ and beyond). However, these methods typically assume conditional independence among the predicted tokens given the prefix, resulting in poor approximations of the true joint distribution over long future spans.
In this work, we propose a different approach: rather than predicting multiple future tokens individually, we train a single auxiliary head to predict a summary of the future sequence. It aims to push as much information as possible about the future into a single target vector, while filtering out information that is inherently unpredictable and would only introduce noise. At each time step $t$, given the future tokens $(x_{t+2},x_{t+3},\dots,x_{t+\tau})$, we construct a summary vector $a(t,\tau)$ to provide supervision to the auxiliary head. We explore two complementary approaches for constructing future summaries: a simple token-level aggregation method, and our main contribution, a learned representation of the future sequence that captures long-range dependencies more effectively.
  • Hand-crafted summary. We train the auxiliary head to predict all the future tokens that will occur in a future window, without requiring to know their exact positions. Concretely, at each step $t$, we define a multi-hot vector $a(t,\tau)$ over the vocabulary, akin to bag-of-words, where $a(t,\tau)_{i}=\mathbb{I}(i\in\{x_{t+2},\dots,x_{t+\tau}\})$, and train the model with a binary cross-entropy objective.
  • Learned summary. Handcrafted summaries such as the one proposed above can be noisy as not all future tokens are equally relevant. To address this, we propose to learn a compact summary of the future. We do this by training a reverse language model ($g_{r}$) on reversed sequences (right-to-left order), so that its hidden representation $a(t,\tau)=g_{r}(x_{\geq t+2})$ serves as a rich embedding of the future context. The auxiliary head of the forward model is then trained to match this representation with an $\ell_{2}$ loss.
Figure 1: A comparison of future-aware pretraining objectives. All methods take a prefix $x_{\leq t}$ as input. NTP: Only predicts the immediate next token. MTP: Uses multiple auxiliary heads, each predicting a specific future token. FSP-BoW: Our proposed hand-crafted summary method that predicts a multi-hot “bag-of-tokens” summary of a long future window using a single auxiliary head. FSP-RevLM: Our proposed learned summary method predicts a compact hidden representation of the future, which is generated by a reverse language model (RevLM).
We use the lens of synthetic tasks, path-star graph and sibling discovery (Bachmann and Nagarajan, 2024; Nagarajan et al., 2025) to clarify the conceptual difference between MTP and our approach. On path-star graph, hand-crafted summaries deliver strong gains over MTP, which only predicts immediate future tokens, demonstrating the benefit of long-horizon supervision. However, in sibling discovery, where the future includes content unrelated to the current local sequence, handcrafted summaries struggle, as they treat all future tokens equally. In contrast, our learned summary vectors focus on the informative parts of the future and achieve consistent gains.
We then scale our approach to real-world pretraining at the 8B parameter level, conducting a systematic evaluation across six methods, including DeepSeek-MTP and multiple handcrafted summary variants. This scale and breadth of comparison is rare in the literature due to its high computational cost. Here, the proposed future summary prediction with learned summaries yields substantial improvements over NTP and MTP baselines, with gains of up to 5% on math and coding benchmarks that demand long-horizon reasoning and planning. These results demonstrate that future summary prediction is not only effective in controlled synthetic settings but also translates into meaningful gains for large-scale LLM training.
2 Future-Aware Pretraining
2.1 Background on Next-token and Multi-token Prediction
Let define a sequence of tokens as $X=(X_{1},X_{2},\dots,X_{T})$ sampled from a distribution $\mathbb{P}_{X}$, where each token is a discrete random variable supported on $\{1,2,\dots,V\}$.
Next-token Prediction (NTP).
The standard next-token prediction objective is:
$L_{\text{NTP}}(X,P_{\theta})=-\mathbb{E}_{x\sim\mathbb{P}_{X}}\Bigg[\sum_{t=1}^{T-1}\log P_{\theta}(x_{t+1}\,|\,x_{\leq t})\Bigg],$ (1)
where $P_{\theta}$ is the predictor, optimized via
$\theta^{\star}=\arg\min_{\theta}L_{\text{NTP}}(X,P_{\theta}).$ (2)
A typical parameterization is:
$P_{\theta}(x_{t+1}\,|\,x_{\leq t})=\mathrm{softmax}\!\Big(f_{u}\circ f_{h}\circ f_{s}\big(x_{\leq t}\big)\Big),$ (3)
where $f_{s}$ is the transformer backbone, $f_{h}$ is a processing head, and $f_{u}$ is the unembedding layer (Figure˜2).
Multi-token Prediction (MTP).
MTP aims to jointly predict multiple future tokens, as follows:
$L_{\text{MTP-Joint}}(X,P_{\theta})=-\mathbb{E}_{x\sim\mathbb{P}_{X}}\Bigg[\sum_{t=1}^{T-1}\log P_{\theta}(x_{t+1},\dots,x_{t+\tau}\,|\,x_{\leq t})\Bigg].$ (4)
Since modeling the exact joint distribution is intractable, a common simplification is to model the marginal distribution of future tokens given the prefix (Gloeckle et al., 2024; Liu et al., 2024):
$L_{\text{MTP}}(X,P_{\theta})=-\mathbb{E}_{x\sim\mathbb{P}_{X}}\Bigg[\sum_{t=1}^{T-1}\sum_{k=1}^{\tau}\mathbf{1}[t+k\leq T]\,\log P_{\theta}(x_{t+k}\,|\,x_{\leq t})\Bigg].$ (5)
Following Gloeckle et al. (2024), the predictor uses separate auxiliary heads for each $k$:
$P_{\theta}(x_{t+1}\,|\,x_{\leq t})$ $=\mathrm{softmax}\!\Big(f_{u}\circ f_{h}\circ f_{s}\big(x_{\leq t}\big)\Big),$ (6)
$P_{\theta}(x_{t+k}\,|\,x_{\leq t})$ $=\mathrm{softmax}\!\Big(f_{u}\circ f^{\prime}_{h_{k}}\circ f_{s}\big(x_{\leq t}\big)\Big),\quad\forall k>1,$
where $f^{\prime}_{h_{k}}$ are auxiliary transformer blocks specialized for predicting future tokens $x_{t+k}$.
This design reduces teacher forcing by predicting $x_{t+k}$ from $x_{\leq t}$ only, rather than conditioning on the full prefix $x_{\leq t+k-1}$. This is the key principle behind multi-token prediction, i.e., reduced teacher forcing by requiring the model to predict a block of future tokens at each step.
We discuss other variants of MTP such as DeepSeek-MTP (Liu et al., 2024), Joint-token Prediction (Ahn et al., 2025), and random future token MTP (Thankaraj et al., 2025) in Appendix 8.
2.2 Future Summary Prediction (FSP)
In MTP, we use a set of auxiliary heads to predict a block of immediate future tokens. A key limitation of this approach is that one does not know exactly where the informative signal in the future sequence lies. Informative signals in future could occur far away in the sequence and be well beyond the number of future tokens ($k$) that MTP predicts. A trivial approach to overcome this could be to predict all the future tokens. However, having one auxiliary head per future token is not scalable, and limits the amount of future tokens we can utilize during training.
Towards this, we propose Future Summary Prediction (FSP), that predicts a compact summary of the (long) future sequence rather than each token individually.
Figure 2: An abstraction of the architecture that subsumes NTP, MTP, and FSP.
Let $a(t,\tau)$ represent a summary of future tokens $(x_{t+2},\dots,x_{t+\tau})$. The learning objective is:
$L_{\text{FSP}}(X,P_{\theta})=L_{\text{NTP}}(X,P_{\theta})+\mathbb{E}_{x\sim\mathbb{P}_{X}}\Big[l_{a}\big(A_{\phi}(x_{\leq t}),a(t,\tau)\big)\Big],$ (7)
where $P_{\theta}$ is the next-token predictor, $A_{\phi}$ is the summary predictor, and $l_{a}$ is the loss between predicted and ground-truth summaries. The architecture is:
$P_{\theta}(x_{t+1}\,|\,x_{\leq t})$ $=\mathrm{softmax}\!\Big(f_{u}\circ f_{h}\circ f_{s}\big(x_{\leq t}\big)\Big),$ (8)
$A_{\phi}(x_{\leq t})$ $=f^{\prime}_{h_{a}}\circ f_{s}\big(x_{\leq t}\big).$
Unlike MTP, FSP requires only a single auxiliary head $f^{\prime}_{h_{a}}$, making it more scalable.
The key question, then, is how to construct an effective future summary. In this work, we investigate two approaches:
(1) Hand-crafted future summaries.
We define a binary vector $a(t,\tau)\in\mathbb{R}^{V}$ indicating whether token $i$ appears in the future window:
$a(t,\tau)_{i}=\mathbb{I}\!\big(i\in\{x_{t+2},\dots,x_{t+\tau}\}\big).$ (9)
Given logits from $A_{\phi}(x_{\leq t})$, we minimize a reweighted binary cross-entropy loss:
$l_{a}\big(A_{\phi}(x_{\leq t}),a(t,\tau)\big)=-\sum_{i=1}^{V}w(i)\Big[a_{i}\log\sigma(z_{i})+(1-a_{i})\log\big(1-\sigma(z_{i})\big)\Big]$ (10)
where $a_{i}=\mathbb{I}\!\big(i\in\{x_{t+2},\dots,x_{t+\tau}\}\big)$, $z_{i}$ is the $i$-th logit of $A_{\phi}(x_{\leq t})$, $\sigma$ is the sigmoid, and $w(i)$ reflects the importance of token $i$ (e.g. term frequency-inverse document frequency, tf–idf).
(2) Learned future summaries.
Predicting every future token via hand-crafted summaries as discussed above could be noisy. For example, not all tokens in the future are informative and thus hand-crafted summaries that account for all of them can be wasteful. Instead of predicting all the future tokens, we therefore propose to predict a learned representation of the future tokens relevant to predicting the current token $x_{t+1}$. We learn this representation via a reverse language model $Q_{\psi}$, (RevLM), trained on "right-to-left" sequences:
$L_{\text{RevLM}}(X,Q_{\psi})=-\mathbb{E}_{x\sim\mathbb{P}_{X}}\Bigg[\sum_{t=1}^{T-1}\log Q_{\psi}(x_{t+1}\,|\,x_{\geq t+2})\Bigg].$ (11)
RevLM shares the same architecture, and we take its hidden state as the summary vector:
$a(t,T-t)=g_{h}\circ g_{s}\big(x_{\geq t+2}\big).$ (12)
We then train $A_{\phi}(x_{\leq t})$ to match this representation via the $\ell_{2}$ loss:
$l_{a}\big(A_{\phi}(x_{\leq t}),a(t,T-t)\big)=\big\|A_{\phi}(x_{\leq t})-g_{h}\circ g_{s}(x_{\geq t+2})\big\|_{2}^{2}.$ (13)
Why Future Summary Prediction?
The key advantage of future summary prediction is its ability to reduce dependence on teacher forcing when modeling long future sequences. To intuitively measure teacher forcing, consider for each ground-truth token exposed to the model, how much information is the model required to predict about unseen tokens? If the model predicts more such information, then we have reduced teacher forcing. Next-token prediction (NTP) uses the highest degree of teacher forcing, since the model always conditions on ground-truth histories to predict just the next token. Multi-token prediction (MTP) partially relaxes this by asking the model to predict short blocks of future tokens, thereby reducing teacher forcing locally. However, MTP remains constrained by the short horizon of its predictions. In contrast, our proposed approach predicts summaries of long future sequences, substantially reducing teacher forcing by requiring the model to reason about rich, global properties of the target trajectory.
3 Analyzing Future Summaries
Future summary prediction provides a unified view of a broad family of pretraining objectives, covering MTP and its variants that sample random future tokens (Thankaraj et al., 2025; Gerontopoulos et al., 2025), and our proposed bag-of-words (FSP-BoW) and ReverseLM (FSP-RevLM) summary prediction objectives.
To better understand how the FSP framework clarifies the tradeoffs of each objective on long-horizon planning, we consider the graph modeling benchmarks introduced in prior works. Our analysis yields following insights:
  • Long future summaries matter. On the canonical path–star task (Bachmann and Nagarajan, 2024), we find that MTP with short range future prediction fails to generalize, highlighting the need for auxiliary objectives that incorporate long-range future information.
  • Adaptive future summaries matter. On a modified sibling discovery task (Nagarajan et al., 2025), we show that incorporating every future token with hand-crafted summaries is suboptimal, highlighting the need for learned summaries that retain key information.
3.1 Long future summary is important
Figure 3: Analysis of FSP-BoW on the path-star task, which tests long-horizon planning. Left: Accuracy (mean over 55 random seeds) of different pretraining objectives on degree 22 graphs with path lengths 66 and 88. Standard NTP generalizes poorly, and MTP’s accuracy degrades as the path length increases, while FSP-BoW achieves perfect accuracy. Right: An illustration of why NTP fails while FSP-BoW succeeds, where the input context is shown in grey and the target is in beige.
We consider the path-star graph, a directed acyclic graph (DAG) $G(d,l)$ composed of $d$ paths, each of length $l$, originating from a central start node $v_{\text{start}}$. The model is given the adjacency list of the graph in the prefix, and the task is to generate the path from $v_{\text{start}}$ to a designated end node $v_{\text{end}}$.
Let the target path be $(v_{\text{start}},v_{1},v_{2},\cdots,v_{\text{end}})$ and the input prefix be $p=(Adj(G),v_{\text{start}},v_{\text{end}})$. Then NTP with teacher forcing predicts an intermediate node in the path as $P_{\theta}(v_{i+1}|p,v_{\leq i})$. As shown by Bachmann and Nagarajan (2024), NTP often learns shortcut solutions: the model can recover $v_{i+1}$ directly from $v_{i}$ by scanning the adjacency list in $p$, without learning the underlying long-range plan (Figure 3). This leads to gradient starvation (Pezeshki et al., 2021), where the supervision signal for the actual planning task is lost. Once the shortcut is learned, meaningful gradient information remains only for predicting the first step $v_{1}$, as it is the only difficult token.
A natural remedy is to reduce teacher forcing via future prediction approaches, which require the model to predict tokens further ahead. This makes shortcuts less effective, as multiple lookups in the adjacency list would be needed to predict future tokens. Hence, we consider the MTP approach, which predicts the immediate future token, and the handcrafted summary method FSP-BoW, which compresses the information about all the future tokens from the path. By summarizing the entire future trajectory, FSP-BoW substantially reduces teacher forcing, encouraging the model to plan the full path instead of exploiting shortcuts.
Findings.
We conduct experiments for two graphs $G(2,6)$ and $G(2,8)$, with all the approaches pretrained from scratch using the GPT-Mini architecture (details in Section˜9.1). At inference, we discard the auxiliary head used in future prediction, and the task is to generate the complete path given the prefix $p$. The evaluation metric checks whether the generated path exactly matches the true path. Results in Figure˜3 (left) show that both NTP and MTP fail to generalize (both obtained perfect training accuracy), and the accuracy of MTP degrades further for the scenario with longer path $G(2,8)$. Hence, predicting just the immediate future token is not enough, and FSP-BoW tackles this by efficiently compressing all the future tokens from the path, enabling it to achieve perfect accuracy in both cases. Appendix 9.1 (Table 4) shows that increasing the number of auxiliary heads in MTP can provide some improvement, but practical limits exist: even with four additional future heads MTP cannot solve the longer path graph $G(2,8)$.
3.2 Adaptive future summary is important
Figure 4: Analysis on the modified sibling discovery task which requires adaptive future summaries. Left: Convergence speed (mean (s.e.) over 33 random seeds) relative to NTP, where lower values imply faster convergence. FSP-RevLM converges faster than NTP while FSP-BoW improves only in the cases with few components. Right: Task setup illustration– given the prefix, only future tokens in the highlighted component are informative. FSP-BoW, which summarizes all future tokens, suffers from irrelevant information, whereas FSP-RevLM summarizes only the informative aspects.
Long handcrafted summaries often incorporate all future tokens, even though only a subset may provide meaningful supervision signals, while the rest can introduce noise. This motivates the need for an adaptive future summary. We illustrate this with modification of the sibling discovery task. Figure 4 (right) depicts the setup, the model must generate sequences made of concatenated, independent components, where each component lists its children nodes first, followed by their parent (e.g., nodes $S_{1}^{1},S_{2}^{1},S_{3}^{1}$ followed by parent $P^{1}$).
Intuitively, the causal factorization implied by the DAG (parent followed by children) implies a goal-conditioned approach: by conditioning on the parent, the model can easily capture the sibling dependencies. Under NTP, the model estimates $P_{\theta}(S_{2}^{1}|S_{1}^{1})$ without the parent, making sibling relationships harder to learn and requiring more samples. Future prediction can help, when the model predicts the parent from $S_{1}^{1}$, then the representation can incorporate the parent information. This enables goal-conditioned planning, the model can predict $S_{2}^{1}$ conditioned on both $S_{1}^{1}$ and the parent, allowing sample efficient learning of sibling relationships (see Nagarajan et al. (2025) for details).
However, not all future tokens are equally informative. As Figure 4 illustrates, future tokens from a different component do not provide relevant signal for predicting $S_{2}^{1}$. Handcrafted summaries that include all future tokens may therefore be affected by the irrelevant information, whereas learned summaries remain robust, as the reverse language model learns representation that emphasize only the predictive signals needed to infer the next token.
Findings.
We verify this empirically by comparing FSP-BoW and FSP-RevLM in experiments with varying number of total components. All models are pretrained from scratch with the GPT-Mini architecture (details in Section˜9.2), and the inference task requires generating a coherent sequence (siblings followed by parent for each component). At convergence, all models produce coherent sequences; to quantify the learning speedup over NTP, we report the ratio of steps to convergence relative to NTP, with lower values indicating faster learning. Since all methods eventually reach perfect consistency at convergence, time to convergence is a reliable metric for comparison. Results in figure 4 (left) shows that FSP-BoW improves over NTP only when the number of components is small, with gains disappearing for more than six components. In contrast, FSP-RevLM consistently achieves faster convergence across all component sizes, confirming the benefit of adaptive future summaries.
4 Experiments
4.1 Setup
We pretrain 3B- and 8B-parameter models on corpora of 250B and 1T tokens, respectively, covering diverse domains. The majority of the data comes from DCLM-like sources and GitHub repositories, supplemented by specialized material in mathematics, programming, and related areas. Models are evaluated across a diverse set of benchmarks: ARC-Easy/Challenge (Clark et al., 2018) for general reasoning, MBPP (Austin et al., 2021) and HumanEval+ (Liu et al., 2023) for code generation, and GSM8K (Cobbe et al., 2021) and Math-500 (Hendrycks et al., 2021) for mathematical reasoning. Details of the pretraining corpus and hyperparameters are provided in Section˜10.1.
We first benchmark our primary method, FSP-RevLM, against the baselines: next-token prediction (NTP), multi-token prediction (MTP) (Gloeckle et al., 2024), and DeepSeek-MTP (DS-MTP) (Liu et al., 2024). Further, a natural way to improve MTP for longer horizons is to add multiple auxiliary heads, each predicting a token farther into the future, but this approach quickly becomes impractical. We therefore constrain both MTP and DS-MTP to a single auxiliary head predicting the immediate future token. This design choice keeps the comparison consistent and aligned with the proposed FSP, since it uses a single auxiliary head.
Building on this unified single auxiliary head framework, we conduct an analysis of how MTP can be enhanced by predicting richer future targets instead of the immediate future token. In addition to predicting the learned future summary (FSP-RevLM), we compare handcrafted future summaries over short- and long-range windows. This includes the proposed bag-of-words future summary (FSP-BoW) as a contributed baseline, and randomly sampled future tokens, in line with prior works (Thankaraj et al., 2025; Gerontopoulos et al., 2025).
Note regarding experiment design. All experiments are conducted under iso-data conditions, meaning that all the methods are trained on identical datasets. For the proposed FSP-RevLM, this implies that both the forward and reverse models are trained on the same data. In line with standard practice in distillation, we do not perform iso-compute comparisons that include the teacher model’s (ReverseLM) cost in the reported compute budget. In practical scenarios, the computational cost of training the teacher model is typically amortized, hence it can be treated as a one-time overhead that is excluded from comparisons of student models (Gemma et al., 2024, 2025).
Note regarding FSP-RevLM implementation. In our experiments with FSP-RevLM, the reverse model is the same size as the forward model (and other baselines), and it is trained for the same number of steps. As a result, FSP-RevLM roughly doubles the total compute cost compared to standard NTP training. While FSP-RevLM increases training FLOPs, we believe this tradeoff is reasonable in today’s compute-rich, data-limited scaling regime (Kim et al., 2025). The field has effectively hit the data wall, whereas available compute continues to grow. Progress increasingly depends on using this growing compute-per-token budget to extract more value from fixed datasets. In this context, methods that deliver measurable gains without requiring additional data, even at higher compute cost, are valuable.
4.2 Results
At the 8B scale (Table 1), future-summary supervision via the ReverseLM (FSP-RevLM) consistently improves the performance across different evaluation tasks. On ARC-Easy, FSP-RevLM ($76.6\%$) provides significant improvement over NTP ($71.8\%$) and MTP ($73.6\%$), and it also leads on ARC-Challenge and MATH. For code generation, it achieves the highest score on MBPP and ties with MTP on HumanEval+, showing that the benefits of predicting future summaries generalize across both reasoning and program synthesis tasks. GSM8K is the one task where NTP ($71.6\%$) holds a lead, though FSP-RevLM ($70.5\%$) still narrows the gap relative to MTP ($67.8\%$).
At the 3B scale (Table 2), DeepSeek-MTP is a strong baseline and obtains better performance than FSP-RevLM on most tasks, except math reasoning. More importantly, FSP-RevLM exhibits larger relative improvements as scale increases from 3B to 8B parameters, and becomes more favourable than DS-MTP. Note that even at the 3B scale, FSP-RevLM still beats MTP on ARC and math reasoning tasks, and performs comparably on MBPP, suggesting that even at smaller scales, learning to predict future summaries may provide more effective auxiliary signal than immediate future token prediction with MTP.
Finally, as a sanity check, in Appendix 10.2 we evaluate these models on standard language modeling task. We find that future-prediction objectives remain competitive with NTP for language modeling, indicating that the reasoning gains over NTP do not come at the cost of degraded language modeling performance.
Task NTP MTP DS-MTP FSP-RevLM ARC-Easy $0.718$ $(0.000)$ $0.736$ $(0.000)$ $0.617$ $(0.003)$ $\bf{0.766}$ $\bf{(0.000)}$ ARC-Challenge $0.531$ $(0.000)$ $0.552$ $(0.000)$ $0.426$ $(0.002)$ $\bm{0.559}$ $(0.000)$ GSM8K $\bf{0.716}$ $\bf{(0.003)}$ $0.678$ $(0.007)$ $0.704$ $(0.003)$ $0.705$ $(0.004)$ MATH $0.342$ $(0.008)$ $0.309$ $(0.006)$ $0.335$ $(0.014)$ $\bm{0.351}$ $\bm{(0.017)}$ MBPP $0.657$ $(0.004)$ $0.672$ $(0.008)$ $0.678$ $(0.006)$ $\bm{0.683}$ $\bm{(0.006)}$ HumanEval+ $0.478$ $(0.019)$ $\bm{0.541}$ $\bm{(0.011)}$ $0.526$ $(0.013)$ $\bm{0.541}$ $\bm{(0.009)}$
Table 1: Pretraining at 8B scale. We benchmark the proposed FSP-RevLM approach against NTP, MTP, and DS-MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and pass@1 for ARC. FSP-RevLM achieves the strongest overall performance, with large gains over the baselines on ARC tasks and MATH, and competitive results with (DS) MTP on code benchmarks.
Task NTP MTP DS-MTP FSP-RevLM ARC-Easy $0.263$ $(0.002)$ $0.272$ $(0.005)$ $\bm{0.293}$ $\bm{(0.000)}$ $0.277$ $(0.000)$ ARC-Challenge $0.263$ $(0.001)$ $0.245$ $(0.002)$ $\bm{0.274}$ $\bm{(0.008)}$ $0.255$ $(0.000)$ GSM8K $0.410$ $(0.003)$ $0.411$ $(0.001)$ $0.417$ $(0.003)$ $\mathbf{0.436}$ $\mathbf{(0.003)}$ MATH $\bm{0.213}$ $\bm{(0.004)}$ $0.196$ $(0.009)$ $0.201$ $(0.004)$ $\bm{0.212}$ $\bm{(0.002)}$ MBPP $0.521$ $(0.007)$ $0.526$ $(0.004)$ $\bm{0.537}$ $\bm{(0.007)}$ $0.524$ $(0.001)$ HumanEval+ $0.301$ $(0.009)$ $0.321$ $(0.015)$ $\mathbf{0.348}$ $\mathbf{(0.022)}$ $0.305$ $(0.006)$
Table 2: Pretraining at 3B scale. We benchmark the proposed FSP-RevLM approach against NTP, MTP, and DS-MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and pass@1 for ARC. At this smaller scale, DS-MTP is a strong overall baseline, but FSP-RevLM outperforms it on math reasoning tasks, and also provides substantial gains over MTP on ARC and math reasoning tasks. Further, as we scale the approaches to 8B parameters, FSP-RevLM scales more favorably than DS-MTP, overtaking it on most tasks.
Method MBPP GSM8K MATH HumanEval+ ARC-Challenge ARC-Easy MTP $0.672$ $(0.008)$ $0.678$ $(0.007)$ $0.309$ $(0.006)$ $\bm{0.541}$ $(0.011)$ $0.552$ $(0.000)$ $0.736$ $(0.000)$ MTP-Skip $\tau$:4 $0.658$ $(0.005)$ $0.639$ $(0.004)$ $0.277$ $(0.020)$ $0.508$ $(0.009)$ $0.494$ $(0.003)$ $0.722$ $(0.000)$ MTP-Skip $\tau$:12 $0.623$ $(0.002)$ $0.621$ $(0.005)$ $0.287$ $(0.018)$ $0.486$ $(0.010)$ $0.512$ $(0.000)$ $0.710$ $(0.003)$ MTP-Skip $\tau$:32 $0.611$ $(0.008)$ $0.598$ $(0.007)$ $0.271$ $(0.005)$ $0.459$ $(0.007)$ $0.379$ $(0.000)$ $0.564$ $(0.000)$ FSP-BoW $\tau$:12 $0.669$ $(0.005)$ $0.699$ $(0.006)$ $0.331$ $(0.016)$ $0.508$ $(0.005)$ $\bm{0.562}$ $\bm{(0.000)}$ $0.737$ $(0.000)$ FSP-BoW $\tau$:100 $0.671$ $(0.002)$ $\bm{0.714}$ $\bm{(0.009)}$ $0.331$ $(0.007)$ $0.500$ $(0.019)$ $0.459$ $(0.000)$ $0.662$ $(0.000)$ FSP-RevLM $0.683$ $(0.006)$ $0.705$ $(0.004)$ $\bm{0.351}$ $\bm{(0.017)}$ $\bm{0.541}$ $(0.009)$ $0.559$ $(0.000)$ $\bm{0.766}$ $\bm{(0.000)}$
Table 3: Analysis of future summary strategies at 8B Scale. We evaluate the effect of different future-summary prediction approaches against vanilla MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and pass@1 for ARC. Handcrafted random-token future summaries (MTP-Skip) perform worse than standard MTP. In contrast, handcrafted multi-hot summaries (FSP-BoW) improve over standard MTP, especially on math reasoning (e.g., GSM8K and MATH), while learned summaries (FSP-RevLM) provide the largest gains across math reasoning and ARC tasks.
Figure 5: Enhanced diversity through learned future summaries. We compare standard MTP, which predicts the immediate future token, with FSP-RevLM, which enriches the auxiliary head target using learned future summaries. FSP-RevLM substantially increases output diversity compared to MTP on GSM8K and MATH benchmarks.
Analyzing MTP with different future summaries.
Table 3 presents our analysis of different future-summary strategies as auxiliary head targets at 8B scale. We focus on the standard MTP architecture, without comparing to DS-MTP as it modifies the input to auxiliary head, to isolate the effect of different future targets on the auxiliary head.
Our results show that random-token handcrafted future summaries (MTP-Skip) perform worse than standard MTP with immediate future token prediction, and performance further degrades as the future window ($\tau$) increases. In contrast, the proposed multi-hot or bag-of-words handcrafted future summaries yield meaningful improvements over MTP, especially on math reasoning tasks, with both shorter $(\tau=12)$ and longer $(\tau=100)$ future windows. For example, FSP-BoW with $\tau=12$ achieves $33.1\%$ on MATH (+2.2 points) and $69.9\%$ on GSM8K (+2.1 points), while even longer windows ($\tau=100$) further amplifies the performance on GSM8K ($71.4\%$, +3.6 points).
Finally, our learned future summaries (FSP-RevLM) outperform MTP across all evaluation tasks, with especially pronounced gains on math reasoning: $35.1\%$ on MATH (+4.2 points) and $70.5\%$ on GSM8K (+3.5 points). Further analysis (Figure 5) for these math reasoning tasks shows that learned summaries promote greater output diversity across different pass@k settings, compared to vanilla MTP with immediate future token prediction. A promising implication of this increased diversity is that post-training on the FSP base model may yield stronger results, as recent work demonstrates that the base model’s solution coverage forms an upper bound on what RL fine-tuning can achieve (Shao et al., 2024; Yue et al., 2025).
In Section˜10.3, we replicate these findings at 3B scale, where both handcrafted and learned future summaries improve over vanilla MTP, again most prominently on math reasoning tasks. Additional ablations explore the effects of omitting reweighting in FSP-BoW and using learned summaries from deeper layers of the ReverseLM teacher model in FSP-RevLM.
(a) 1T Unique Tokens (1 Epoch)
(b) 50B Unique Tokens (20 Epochs)
Figure 6: Data-constrained pretraining at 8B scale. Left: standard pretraining with 1T unique tokens for 1 epoch (same as Table˜1). Right: data-constrained pretraining with 50B unique tokens for 20 epochs, matched for total compute. We report average accuracy across all reasoning tasks. FSP-RevLM outperforms NTP in both regimes, while in the data-constrained setting all future-aware objectives surpass NTP, indicating that richer future-aware supervision is especially useful when unique data are limited.
Data-constrained Pretraining Results.
Motivated by recent work highlighting data-constrained pretraining regimes, in which repeated exposure to a limited pool of unique tokens becomes increasingly common (Muennighoff et al., 2023; Kim et al., 2025), we also evaluate our 8B models in this setting. Specifically, we train on 50B unique tokens for 20 epochs, so that the total token budget matches the 1T-token budget used in the 1-epoch setup considered previously in Table˜1. This enables an iso-compute comparison of pretraining objectives as the supply of unique data is reduced. Figure 6 reports average performance across all reasoning tasks, with the left sub-figure corresponding to the standard 1-epoch regime and the right sub-figure to the data-constrained 20-epoch regime. We find that FSP-RevLM continues to outperform NTP in the data-constrained setting, while remaining competitive with MTP variants. Notably, in the data-constrained regime all future-aware objectives outperform NTP, including DS-MTP, which underperformed in the 1-epoch setting. Overall, these results reinforce the view that when the total unique tokens are limited, richer future-aware supervision can extract more useful learning signal from each token than next-token prediction alone. Further, in Appendix 10.4 we present a more comprehensive analysis at the 3B scale, showing that as the budget of unique tokens decreases, future-aware methods outperform NTP even on standard language modeling metrics.
5 Related Work
Beyond Multi-token Prediction. Multi-token Prediction (MTP) extends Next-token Prediction (NTP) by introducing auxiliary heads to predict future tokens (Gloeckle et al., 2024), thereby providing a richer learning signal. Variants such as DeepSeek-MTP (Liu et al., 2024), Joint-token Prediction (Ahn et al., 2025), and Next-latent Prediction (Teoh et al., 2025) differ in how they structure these auxiliary predictions, but they retain a common limitation: scaling auxiliary heads to capture long-range dependencies becomes impractical. More recent approaches attempt to improve efficiency by predicting randomly selected or non-sequential future tokens (Gerontopoulos et al., 2025; Thankaraj et al., 2025; Liu et al., 2025). However, heuristically sampling future tokens risks missing the most informative long-range signals. Our approach (FSP-RevLM) addresses this by predicting a learned future summary, such that it can extract meaningful long-range information.
Related to the proposed hand-crafted summary approach (FSP-BoW), Yin et al. (2024) and Frydenlund (2025) incorporate objectives that predict an unordered set-level summary of future tokens, conceptually similar to our FSP-BoW and its binary cross-entropy formulation. However, our formulation is more general as we allow re-weighting schemes such as tf-idf scores within the binary cross-entropy objective and is implemented via auxiliary heads to be consistent with the multi-token prediction architecture (Gloeckle et al., 2024). More importantly, our contribution is not limited to a specific BoW variant: we introduce a unified framework for future summary prediction that clarifies the trade-offs between different forms of future supervision. For instance, a BoW summary directly resolves the path–star graph failure by aggregating the long-range information, yet it remains suboptimal when only part of the future is relevant, as observed in our sibling discovery experiments. These insights motivate moving beyond hand-crafted summaries toward a learned future summary via reverse LM, which adaptively captures the most informative aspects of the future.
Finally, regarding SemFormer (Yin et al., 2024), beyond its BoW-style supervision, its central idea is to introduce special planning tokens that are trained to predict a latent embedding of the future sequence, encouraging planning. While conceptually related, it applies future supervision only at designated planning tokens and relies on auto-encoding objective to learn future embeddings. In contrast, our approach requires no special planning tokens, enforces future-summary alignment at every position in the sequence through auxiliary heads, and learns future embeddings via the ReverseLM, yielding a fundamentally different supervision signal.
Leveraging the right-to-left signal.
Leveraging reverse, or "right-to-left", order during training has shown empirical benefits across multiple learning paradigms. The Belief State Transformer (BST) (Hu et al., 2024) employs dual forward and backward encoders to predict both the next token after a prefix and the previous token before a suffix. This bidirectional training encourages the model to form a compact belief state, though it does not explicitly address teacher forcing. In contrast, our FSP-RevLM incorporates a reverse model with a different objective: mitigating dependence on teacher forcing. While trained on standard left-to-right sequences, FSP-RevLM aligns the forward model’s embeddings with those of a reverse model, effectively distilling the reverse order into the forward language model.
The closest prior work on reducing teacher forcing is Twin Networks (Serdyuk et al., 2017), which trains a reverse RNN to align forward hidden states to those of the reverse model to encourage long-range future dependence. While similar at a high level, our contributions go beyond the specific mechanism of generating future summaries via reverse language models. We cast future summary prediction as a broader framework for pretraining objective design, and provide evidence for when simple handcrafted summaries fail and when learned summaries from a reverse LM are needed. Moreover, extending this idea for large-scale pretraining with transformers is itself non-trivial. Just as Twin Networks anticipated aspects of framework, works predating modern MTP (Gloeckle et al., 2024) had already considered predicting multiple future tokens or parallel token blocks from a prefix (Tschannen et al., 2023; Monea et al., 2023), but did not establish such objectives as effective for large-scale transformer pretraining. Gloeckle et al. (2024) made that step for modern LLMs, and our work extends it by showing that FSP can be scaled for better pretraining.
Finally, Meet-in-the-Middle (MiM) (Nguyen et al., 2023) jointly trains forward and reverse models with shared parameters and aligns their output distributions as a regularizer. Our method differs in two key aspects: (1) we perform distillation with the reverse model rather than parameter sharing, and (2) we avoid the somewhat impractical inductive bias of MiM that the forward and reverse output distributions must match exactly.
6 Conclusion
In this work, we highlighted a key limitation of existing multi-token prediction methods: the difficulty of scaling auxiliary heads for long-horizon future prediction. Towards this, we proposed Future Summary Prediction (FSP), a novel pretraining framework that shifts the auxiliary objective from predicting specific future tokens to predicting a (learned) summary of the future. Experiments on 8B models, using both hand-crafted and learned summaries, demonstrate that FSP delivers a stronger, more robust training signal, improving over NTP and MTP on challenging reasoning and coding tasks. Our findings indicate that focusing on abstract, predictable aspects of the future is a promising strategy for designing more efficient and effective pretraining objectives for next-generation large language models. Future work can focus on designing computationally efficient approaches for learning future summaries (embeddings), potentially without training a reverse model.
Acknowledgments
The authors thank Sachin Mehta for suggestions on the pretraining experiment setup, and Vaishnavh Nagarajan and Nanda H Krishna for their detailed feedback on the draft. The authors also gratefully acknowledge helpful discussions with Aniket Didolkar, Andrei Nicolicioiu, Mathurin Videau, Sarthak Mittal, Sharut Gupta, Vineet Jain, and Moksh Jain. This research was fully funded by Meta under the AI Mentorship (AIM) program with Mila. Ioannis Mitliagkas in his role as Divyat Mahajan’s academic advisor, acknowledges support by an NSERC Discovery grant (RGPIN-2019-06512), and Canada CIFAR AI chair. Divyat Mahajan acknowledges support via FRQNT doctoral scholarship (https://doi.org/10.69777/354785) for his graduate studies.